# Before & After: The Effect of EU’s 2022 Code of Practice on Disinformation

### v1

@@ Abstract.
In this work, we set out to explore what was the impact of the Code of Practice and especially to explore to what extent ad networks continue to advertise on dis-/mis-information sites. We perform a historical analysis and find that, although at a hasty glance things may seem to be improving, there is really no significant reduction in the amount of advertising relationships among popular misinformation websites and major ad networks. In fact, we show that ad networks have withdrawn mostly from unpopular misinformation websites with very few visitors, but still form relationships with highly unreliable websites that account for the majority of misinformation traffic. To make matters worse, we show that ad networks continue to place advertisements of legitimate companies next to misinformation content. In fact, major ad networks place ads in almost 400 misinformation websites of our dataset.
Disinformation, Online Advertising, Code of Practice on Disinformation, ads.txt, sellers.json
ccs: Information systems World Wide Webccs: Information systems Online advertising
@@ 1. Introduction
From the US presidential elections in 2016 (bovet2019influence, ) to the Brexit referendum (marshall2018post, ), Covid-19 (rocha2021impact, ), the war in Ukraine (khaldarova2020fake, ) and the recent conflict in the Middle East (middleeast, ), disinformation today is shaping major global events. It tears apart the fabric that holds our societies together by destroying people’s faith in traditional news sources and undermining people’s trust in governments, public institutions and democratic processes (ognyanova2020misinformation, ). Disinformation is usually designed to appeal to our worst impulses, fears and prejudices – all in an attempt to shape opinions or divide and conquer a society on the information battlefield. According to a recent survey, over 80% of EU citizens see fake news as an issue both for their country, and for democracy in general (fakenews_survey2018, ).
In an attempt to mitigate this threat, governments, tech firms and stakeholders have explored various methods to identify and curtail the spread of fake news. In 2022, the United Nations in their report (un_report, ) set out the relevant international legal framework, and discussed measures that States and technology enterprises reported to have taken to counter disinformation. In 2022, Google launched the Google’s Digital News Innovation Fund that supported 662 digital news projects in Europe (total amount: 150M euros) (dnifund, ). A few years earlier, the European Union issued the 2018 Code of Practice on Disinformation, where representatives of online platforms, leading tech companies and players in the advertising industry (e.g., Facebook, Twitter, Google, Mozilla with Microsoft and TikTok following) agreed on a self-regulatory Code of Practice to address the spread of online disinformation (cop2018, ). The Code of Practice is based on 21 commitments in different domains, including the transparency in political advertising and, more importantly, demonetization of purveyors of disinformation.
A year after its implementation, the Code was already acknowledged globally as a pioneering framework (covidAssessment, ). An assessment of the Code was published in 2020 (assessment, ), showing that (i) the Code did provide a valuable framework for a structured dialogue between online platforms, thus, ensuring transparency and accountability of their policies on disinformation, but also (ii) a set of important gaps and shortcomings. In June 2022, the Commission issued the Strengthened Code of Practice on Disinformation (CoP) (codeOfPractice, ).
In this paper, we set out to explore the impact of the Strengthened CoP with regards to one of its most important commitments: the demonetization of disinformation spreading websites. We perform a historical analysis on misinformation websites in order to investigate what was the impact of the policy on their ad revenue by comparing their direct business relationships with ad networks.
@@
    We discover that there is no significant reduction in the amount of advertising accounts with popular ad networks, after the Code of Practice on Disinformation came into effect.
  2. (2)
    We demonstrate that after the CoP came into effect, popular ad networks have withdrawn mostly from unpopular disinformation websites, with hardly any visitors. On the contrary, we find that these ad networks still have connections with popular misinformation websites that account for the majority of misinformation traffic.
@@
    We show that ad networks do not substantially differentiate between unreliable and reliable news websites in their business relations and that they form business relationships with both kinds of websites.
@@
    We make our list of approximately 2,500 misinformation news websites publicly available (openSourceData, ).
@@ 2. Background
In the modern Web, when users visit a website, an automated programmatic process takes place to match advertisers with users. Advertisers entrust Demand-Side Platforms (DSPs) to find the most suitable users to promote their brand or product, while website publishers authorize Supply-Side Platforms (SSPs) to manage their ad inventory and monetize their content. This operation consists of multiple entities, numerous intermediaries and complex interactions. To eradicate various forms of ad fraud, stakeholders have introduced new standards.
2.1. ads.txt
@@
Due to the complexity of the advertising ecosystem, bad actors were able to sell the ad inventory of websites they did not control (i.e., counterfeit inventory). In an attempt to eliminate this type of ad fraud, the the Internet Advertising Bureau (IAB) Technology Laboratory introduced the ads.txt specification (adsTxtStandard, ). Using this standard, websites publishers explicitly disclose who is authorized to sell the ad inventory of a website. This information is enclosed within a text file which as publicly available at the root of the domain. We provide a snipped of an ads.txt file in Figure 1. Each ads.txt file contains one or many comma-separated records with each record authorizing an entity to sell the ad inventory of the website. Each record contains: (i) the domain name of the advertising system (e.g., SSP or Ad Exchange), (ii) an identifier for the seller’s account in the advertising system, (iii) the type of the account, and optionally (iv) a certification id for the advertising system. According to the specification (adsTxtStandard, ), the type of the account in field (iii) represents the relationship between the account holder and the advertising system. A DIRECT account suggests that the website publisher also controls the advertising account and has a direct business contract with the ad system. A RESELLER accounts indicates that the website publisher has authorized a third-party entity to resell their ad inventory.
2.2. sellers.json
@@
To increase transparency in the advertising ecosystem, the IAB Tech Lab also introduced the sellers.json standard (sellersJsonStandard, ). This standard helps discover the identities of the entities that sell the ad inventory of websites and can help identify all intermediaries involved in ad inventory selling. The sellers.json standard is implemented as a JSON file published by advertising systems in order to disclose the inventory sellers they have approved within their ad system. Each approved seller is represented by a unique identifier inside this file. This is the same identifier that appears in the ads.txt file and maps to a single entity that is paid for the ad inventory it sells. We provide an example in Figure 1. Additionally, each identifier is described by a seller type. A PUBLISHER seller suggests that this account directly owns and controls the website whose ad inventory is sold through the advertising system and that the advertising system directly pays this entity. An INTERMEDIARY sellers indicates that is a reseller of ad inventory. Finally, a seller can be labeled as BOTH, acting as both a publisher and intermediary.
@@
Figure 1. Snippets of the ads.txt file served by Tumblr and the sellers.json file served by Google on April 2024. The ID that describes an advertising account matches in both files.
@@ 3.1. Sources of Misinformation
To assemble a list of misinformation websites and study their behavior over time, we make use of publicly available datasets produced by academic publications, as well as by credible and acclaimed organizations. Our first sources is MediaBias/FactCheck (MBFC) (mbfc, ), an organization that investigates bias and misinformation in online sources of information. By using human evaluators and a concrete and consistent methodology (mbfcMethodology, ), they evaluate the factual accuracy of various media sources. We visit the MBFC website in October 2023 and extract labels and information about the bias, factual reporting, credibility, traffic volume and organization type for each evaluated website. We retain only websites that have a “LOW CREDIBILITY” score and a factual reporting of “MIXED”, “LOW” or “VERY LOW” that form a list of 1,677 misinformation websites. We consciously exclude websites from the “Satire” category from this analysis, since they deliberately use humor and exaggeration in their published articles and do not attempt to deceive visitors. Such websites do not fall under the misinformation umbrella.
Next, we use the publicly available lists of previous academic works to extend our dataset of fake news websites as assembled in (papadogiannakis2023funds, ) consisting of 1,044 misinformation websites. Additionally, we extract 669 more misinformation sites from (vekaria2023inventory, ) that classifies news websites using independent sources (e.g., Politifact, Snopes). Finally, we make use of the FakeNewsNet tool used in (shu2020fakenewsnet, ) and download all news articles from the collected dataset. We exclude all data from the GossipCop fact-checking service because it has been discontinued and is no longer available. We extract 8 misinformation websites by filtering websites that have published at least 4 fake news articles (proven by Politifact (politifact, )). We also filter out Facebook because we argue that it is neither a news publisher or news aggregator. Misinformation articles or posts shared in such a big platform should not classify the entire social network as a misinformation website.
We merge all the above lists and end up with a dataset of 2,469 distinct misinformation websites. In an attempt to foster academic research and give back to the academic community, we make our list of misinformation websites publicly available (openSourceData, ).
@@
A big portion of our study revolves around the analysis of ads.txt and sellers.json, two important files in the digital advertising that help combat ad fraud and increase transparency in the advertising ecosystem. Previous work has demonstrated that these files can reveal business relationships among websites and ad networks (bashir2019longitudinal, ; papadogiannakis2023funds, ; vekaria2023inventory, ). We first study the business relationships among ad networks and misinformation websites and how these have changed over the years and especially after the Code of Practice on Disinformation was signed. To that extent, we utilize ads.txt served by misinformation websites, and the sellers.json files served by popular ad networks. We crawl the list of approximately 2,500 misinformation websites on January 2024 and download 226,659 ads.txt entries coming from 1,132 misinformation websites. We make use of an open-source crawler (papadogiannakis2023pooling, ) to collect and parse the respective ads.txt files. We find there are ads.txt records representing relationships with over 1,100 distinct ad networks, suggesting that ad networks still work with objectionable websites.
To get a better understanding of the actual business relationships, among misinformation websites and ad networks, in this work we direct our analysis to ads.txt entries marked as DIRECT. The ads.txt specification defines that DIRECT accounts indicate the website owner (i.e., publisher) directly controlling the advertising account, and that there is a direct business contract between the publisher and the ad system (adsTxtStandard, ). Note that a website may provide bogus information in its ads.txt file. For example, it may claim that a legitimate advertiser sells ads on this web site. On the other hand, an ad seller may also provide a fake sellers.json file claiming that it can put ads in high-profile web sites in order to attract more customers.
Although each individual file of the above two files may provide fake information, their intersection provides the truth. Indeed, if website A (in its ads.txt file) claims that it displays content from ad network B, and ad network B (in its sellers.json file) claims that it displays ads in website A, then this is true. That is why in the rest of the paper we take the intersection of these files. That is, we form a relation between A and B only if we find this relation both in ads.txt file of A and in sellers.json file of B.
@@ 3.3. Ad Collection
First, we establish different user profiles (i.e., personas) to emulate real users and present a compelling profile to ad networks. We form different personas emulating different real-world users so that they receive ads from different advertisers, thus revealing a better picture of the ad ecosystem around misinformation. Additionally, we emulate users from various geographic locations in order to (potentially) get ads from both acclaimed companies and brands known worldwide, as well as local advertisers. We select Netherlands, Greece and USA as the different vantage points. We use a VPN service to set the location of our virtual users and before we begin forming the personas, we use two different geolocation services to ensure that these virtual users indeed appear to be in the specified country.
In order to build a persona, we visit a set of specific websites and expect that third parties will infer the user’s gender and preferences based on the content they consume. To form a list of websites that are associated with a gender of a specific country, we utilize the SimilarWeb analytics platform (similarWeb, ). Specifically, we visit SimilarWeb and extract the 5 most popular websites in each category and each country of our experiment. SimilarWeb uses a taxonomy system with 210 distinct website categories (including subcategories). For each website, we extract its demographic data (if available). SimilarWeb uses two gender groups (i.e., female and male) to classify visitors of websites. Next, we sort websites based on their audience’s gender representation and ensure that all websites have a representation of at least 55% for that gender. We set an upper threshold of 200 websites and a lower threshold of 100 websites to train each persona while all. The lower threshold is more than enough since previous work has demonstrated that a persona is stable after visiting the top 10 websites associated with users of that demographic (agarwal2020stop, ). This process leads to 6 (combination of 2 genders and 3 countries) lists of website that represent the interests of the specific persona.
We utilize Playwright (playwright, ), a browser automation framework and set up legitimate browser profiles, retaining cookies and local storage. To build the user profiles, we try to emulate the behavior of a real user that visits some websites to consume online content. We try to make the behavior of our automated crawler as non-deterministic as possible to appear as a legitimate user. The behavior we are emulating is that of a user that visits a website, reads its content and then navigates through the website after quickly inspecting the content of each page. For each website in our target list we (i) visit the landing page, (ii) wait for a random number of seconds in range [2, 5], and (iii) explore 5 subpages. In order to explore a subpage our crawler clicks on a randomly selected internal URL and navigates to this subpage, waits for a random number of seconds to emulate a user consuming content, scrolls for a random number of seconds and then waits for an additional random number of seconds. The websites of the list are visited in a random order and after all websites have been visited, the user profile is stored.
@@
After we have built each persona, we use this user profile to visit the list of misinformation websites. We utilize the open-source crawler presented in (papadogiannakis2022leveraging, ) and extend it to cover the needs of this work. The crawler visits both the landing page and a randomly selected internal page of each website and detects the advertisers in these. We make this decision based on previous work that has demonstrated that internal pages can have significant differences in both structure and content (including ads and trackers) (10.1145/3419394.3423626, ). We extend the crawler to also detect online ads using cosmetic filters, based on Easylist  [^1] https://easylist.to/ and uBlock Origin  [^2] https://ublockorigin.com/, two of the most popular ad-blocking projects. Additionally, to make the crawler appear as close to an actual user as possible, we use a headful browser with custom preferences (e.g. dark color scheme, window size, window position). Finally, we make use of Consent-O-Matic (nouwens2022consent, ), a reliable extension, to automatically accept all cookies in consent banners and ensure that ads are not blocked because the user has not interacted with the consent banner. We configure the extension to automatically accept all data processing purposes before we visit any website.

### v2

@@ Abstract.
In this work, we set out to explore what was the impact of the Code of Practice and especially to explore to what extent ad networks continue to advertise on dis-/mis-information sites. We perform a historical analysis and find that, although at a hasty glance things may seem to be improving, there is really no significant reduction in the amount of advertising relationships among popular misinformation websites and major ad networks. In fact, we show that ad networks have withdrawn mostly from unpopular misinformation websites with very few visitors, but still form relationships with highly unreliable websites that account for the majority of misinformation traffic. To make matters worse, we show that ad networks continue to place advertisements of legitimate companies next to misinformation content. We show that major ad networks place ads in almost 400 misinformation websites in our dataset.
Misinformation; Disinformation; Advertising; Web Monetization; Code of Practice
journalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714898isbn: 979-8-4007-1274-6/25/04ccs: Information systems World Wide Webccs: Information systems Online advertising
@@ 1. Introduction
From the US presidential elections in 2016 (Bovet and Makse, 2019) to the Brexit referendum (Marshall and Drieschova, 2018), Covid-19 (Rocha et al., 2021), the war in Ukraine (Khaldarova and Pantti, 2020) and the recent conflict in the Middle East (CHIM et al., 2023), disinformation today is shaping major global events. It tears apart the fabric that holds our societies together by destroying people’s faith in traditional news sources and undermining people’s trust in governments, public institutions, and democratic processes (Ognyanova et al., 2020). Disinformation is usually designed to appeal to our worst impulses, fears and prejudices – all in an attempt to shape opinions or divide and conquer a society on the information battlefield. Over 80% of EU citizens see fake news as an issue both for their country, and for democracy in general (Sandford, 2018).
In an attempt to mitigate this threat, governments, tech firms, and stakeholders have explored various methods to identify and curtail the spread of fake news. In 2018, Google launched the Digital News Innovation Fund that supports over 7000 news partners (\$300M in funding) (Initiative, 2022). In 2022, the United Nations in their report (Secretary-General of the United Nations, 2022) set out the relevant international legal framework, and discussed measures that States and technology enterprises reported to have taken to counter disinformation. A few years earlier, the European Union issued the 2018 Code of Practice on Disinformation, where representatives of online platforms, leading tech companies and players in the advertising industry (e.g., Facebook, Twitter, Google with Microsoft and TikTok following) agreed on a self-regulatory Code of Practice to address the spread of online disinformation (Commission, 2018). The Code of Practice is based on 21 commitments in different domains, including the demonetization of purveyors of disinformation.
A year after its implementation, the Code was already acknowledged globally as a pioneering framework (Commission, 2020b). An assessment of the Code was published in 2020 (Commission, 2020a), showing that (i) the Code did provide a valuable framework for a structured dialogue between online platforms, thus, ensuring transparency and accountability of their policies on disinformation, but also (ii) a set of important gaps and shortcomings. In June 2022, the Commission issued the Strengthened Code of Practice on Disinformation (CoP) (Comission, 2022a).
In this paper, we set out to explore the impact of the Strengthened CoP with regard to one of its most important commitments: the demonetization of disinformation-spreading websites. We perform a historical analysis on misinformation websites in order to investigate what was the impact of the policy on their ad revenue by comparing their direct business relationships with ad networks.
@@
    We discover that there is no significant reduction in the amount of advertising accounts with popular ad networks, after the Code of Practice on Disinformation came into effect. Popular ad networks have withdrawn mostly from unpopular disinformation websites, with hardly any visitors. On the contrary, we find that these ad networks still have connections with popular misinformation websites that account for the majority of misinformation traffic.
@@
    We show that ad networks do not substantially differentiate between unreliable and reliable news websites in their business relations, and that they form business relationships with both kinds of websites.
@@
    We make our list of approximately 2,500 misinformation news websites publicly available (Papadogiannakis, 2024).
@@ 2. Background
When users visit a website, an automated programmatic process matches them with advertisers. Advertisers rely on Demand-Side Platforms (DSPs) to target the right users, while website publishers use Supply-Side Platforms (SSPs) to manage their ad inventory.
@@
ads.txt Due to the complexity of the advertising ecosystem, bad actors were able to sell ad inventory of websites they did not control. To eliminate this, the Internet Advertising Bureau (IAB) Technology Laboratory introduced the ads.txt specification (Laboratory, 2022) that allows publishers to explicitly disclose who is authorized to sell the ad inventory of their website. This information is enclosed within a text file, publicly available at the root of the domain (snippet in Figure 3). Each ads.txt file contains comma-separated records, with each record authorizing an entity to sell the ad inventory of the website. Each record contains: (i) the domain name of the advertising system (e.g., SSP), (ii) an identifier for the seller’s account, (iii) the type of the account, and optionally (iv) a certification ID for the advertising system. The type of the account represents the relationship between the account holder and the ad system. A DIRECT account suggests that the website publisher also controls the advertising account and has a direct business contract with the ad system. For a RESELLER account, the website publisher has authorized a third-party entity to resell their ad inventory.
@@
sellers.json To increase transparency in the advertising ecosystem, the IAB Tech Lab also introduced the sellers.json standard (Laboratory, 2019). This standard helps discover the identities of the entities that sell the ad inventory of websites and can help identify all intermediaries involved in ad inventory selling. The sellers.json standard is implemented as a JSON file published by advertising systems in order to disclose the inventory sellers they have approved within their ad system. Each approved seller is represented by a unique identifier inside this file. This is the same identifier that appears in the ads.txt file and maps to a single entity that is paid for the ad inventory it sells (example in Figure 3). Additionally, a seller type describes each identifier. A PUBLISHER seller suggests that this account directly owns and controls the website whose ad inventory is sold through the advertising system, and that the advertising system directly pays this entity. An INTERMEDIARY record indicates that it is a reseller of ad inventory. Finally, a seller can be labelled as BOTH, acting as both a publisher and an intermediary.
@@
Figure 1. Snippets of Tumblr’s ads.txt file and Google’s sellers.json file on April 2024. The ID that describes an advertising account matches in both files.
@@ 3.1. Sources of Misinformation
Disinformation is the deliberate creation or spread of false information, whereas misinformation is false information shared without necessarily intending to deceive. While we study the Code of Practice on Disinformation, our focus is on how misinformation websites generate ad revenue. Proving that there is clear intent to mislead (i.e., disinformation) is not a trivial task and falls outside the scope of this work. Nonetheless, some websites in our list have been labeled as disinformation spreaders (e.g., (Check, 2024e, d)). We acknowledge the distinction between these two concepts but argue that understanding the business model behind false content is crucial, as many disinformation campaigns rely on the same advertising mechanisms as misinformation (Silverman et al., 2022; Skibinski, 2021). The ability to profit from misleading content encourages bad actors to deliberately spread false claims, blurring the line between misinformation and disinformation.
To assemble a list of misinformation websites and study their behavior over time, we make use of publicly available datasets produced by academic publications, as well as by credible and acclaimed organizations. Our first source is MediaBias/FactCheck (MBFC) (Check, 2023), an organization that investigates bias and misinformation in online sources of information. By using human evaluators and a concrete and consistent methodology (Check, 2024b), they evaluate the factual accuracy of various media sources. We visit the MBFC website in October 2023 and extract labels and information about the bias, factual reporting, credibility and traffic volume for each evaluated website. We retain only websites that have a “LOW CREDIBILITY” score and a factual reporting of “MIXED”, “LOW” or “VERY LOW”, forming a list of 1,677 misinformation websites. We consciously exclude websites from the “Satire” category from this analysis, since they deliberately use humor and exaggeration in their published articles and do not attempt to deceive visitors. Such websites do not fall under the misinformation umbrella.
Next, we use the publicly available lists of previous academic works. We extend our dataset with fake news websites assembled in (Papadogiannakis et al., 2023) consisting of 1,044 misinformation websites. Additionally, we extract 669 more misinformation sites from (Vekaria et al., 2024a) that classifies news websites using independent sources (e.g., Politifact, Snopes). Finally, we make use of the FakeNewsNet tool used in (Shu et al., 2020) and download all news articles from the collected dataset. We exclude all data from the GossipCop fact-checking service because it has been discontinued and is no longer available. We extract 8 misinformation websites by filtering websites that have published at least 4 fake news articles (proven by Politifact (Institute, 2023)). We also filter out Facebook because we argue that it is neither a news publisher nor a news aggregator. Misinformation posts on a large platform shouldn’t label the entire network as misinformation.
We merge all the above lists and end up with a dataset of 2,469 distinct misinformation websites. In an attempt to foster academic research and give back to the academic community, we make our list of misinformation websites publicly available (Papadogiannakis, 2024).
@@
Our study revolves around ads.txt and sellers.json, two important files in digital advertising that help combat ad fraud and increase transparency in the advertising ecosystem. Previous work has demonstrated that these files can reveal business relationships among websites and ad networks (Bashir et al., 2019; Papadogiannakis et al., 2023; Vekaria et al., 2024a). We first study the business relationships among ad networks and misinformation websites and how these have changed over the years and especially after the Code of Practice on Disinformation was signed. To that extent, we utilize ads.txt files served by misinformation websites, and the sellers.json files served by popular ad networks. We crawl the list of approximately 2,500 misinformation websites on January 2024 and download 226,659 ads.txt entries coming from 1,132 misinformation websites (an ads.txt file of a single website can contain multiple entries for multiple ad networks (Papadogiannakis et al., 2025)). We make use of an open-source crawler (Papadogiannakis et al., 2025) to collect and parse the respective ads.txt files. We find there are ads.txt records representing relationships with over 1,100 distinct ad networks, suggesting that ad networks still work with objectionable websites. To get a better understanding of the actual business relationships, we direct our analysis to ads.txt entries marked as DIRECT. The ads.txt specification defines that DIRECT accounts indicate the website owner (i.e., publisher) directly controls the advertising account, and that there is a direct business contract between the publisher and the ad system (Laboratory, 2022). Note that a website may provide bogus information in its ads.txt file. For example, it may claim that a legitimate advertiser sells ads on this website. On the other hand, an ad seller may also provide a fake sellers.json file claiming that it can put ads in high-profile websites to attract more customers. Although each individual file of the above two files may provide fake information, their intersection provides the truth. Indeed, if website A (in its ads.txt file) claims that it displays content from ad network B, and ad network B (in its sellers.json file) claims that it displays ads in website A, then this is true. We provide an example in Figure 3. In the rest of this work we take the intersection of these files. That is, we form a relation between A and B only if we find this relation both in ads.txt file of A and in sellers.json file of B.
@@ 3.3. Ad Collection
First, we establish different user profiles (i.e., personas) to emulate real users and present a compelling profile to ad networks. We form different personas emulating different real-world users so that they receive ads from different advertisers, thus revealing a better picture of the ad ecosystem around misinformation. Additionally, we emulate users from various geographic locations in order to (potentially) get ads from both acclaimed companies and brands known worldwide, as well as local advertisers. We select Netherlands, Greece, and USA as the different vantage points. We choose two EU countries since the CoP is an EU Commission initiative. Given its self-regulatory nature and potential differences outside the EU, we also include the US to identify ad servings that may not be evident in the EU. We use a VPN service to set the location of our virtual users, and before we begin forming the personas, we use two different geolocation services to ensure that these virtual users indeed appear to be in the specified country. In order to build a persona, we visit a set of specific websites and expect that third parties will infer the user’s gender and preferences based on the content they consume. We describe the methodology we follow to build user profiles in Appendix A.
Figure 5. Direct business relationships of top ad networks with unpopular misinformation websites after 2 years.
@@
After we have built each persona, we use this user profile to visit the list of misinformation websites. We utilize the open-source crawler presented in (Papadogiannakis et al., 2022) and extend it to cover the needs of this work. The crawler visits both the landing page and a randomly selected internal page of each website and detects the advertisers in these. Prior work has demonstrated that internal pages can have significant differences in both structure and content (including ads and trackers) (Aqeel et al., 2020). We extend the crawler to also detect online ads using cosmetic filters, based on Easylist  [^1] https://easylist.to/ and uBlock Origin  [^2] https://ublockorigin.com/, two of the most popular ad-blocking projects. Additionally, to make the crawler appear as close to an actual user as possible, we use a headful browser with custom preferences (e.g., dark color scheme, window size, etc.). Finally, we make use of Consent-O-Matic (Nouwens et al., 2022), a reliable extension, to automatically accept all cookies in consent banners and ensure that ads are not blocked because the user has not interacted with the consent banner. We configure the extension to automatically accept all data processing purposes.